OpenStudy (anonymous):

solve 3 log 2x=4 round to the nearest thousandth if necessary

OpenStudy (anonymous):

log2x = 4/3 10^log2x = 10^4/3 2x = 10^4/3

OpenStudy (michele_laino):

hint: \[\large 3\log \left( {2x} \right) = \log \left\{ {{{\left( {2x} \right)}^3}} \right\}\]